# ConfTuner: Training Large Language Models to Express Their Confidence Verbally

### v1

@@ 1 Introduction
A large language model’s (LLM) ability to recognize and communicate uncertainty through verbalized confidence–that is, expressions of confidence conveyed in natural language, such as “I am 80 percent confident that…” (teaching, )–is central to effective human–AI collaboration (li2025as, ). This capability is particularly important in high-stakes domains such as scientific inquiry (asai2024openscholar, ), law (li2024legal, ), and healthcare (li2024mediq, ), where decision quality and interpretability are essential. However, current LLMs are not explicitly trained to express calibrated uncertainty. As a result, they often report very high confidence even when producing hallucinated or incorrect content (hallucination1, ; hallucination2, ; hallucination3, ; canllm, ). This overconfidence problem undermines trust and poses serious challenges for the safe deployment of LLMs (Figure 1).
Recent efforts (just, ; canllm, ; teaching, ; sayself, ; lacie, ) have focused on improving the elicitation of verbalized confidence from LLMs. Prompt-based methods rely on carefully crafted instructions (just, ; canllm, ), but have shown limited effects in improving calibration (just, ; canllm, ). Alternatively, training-based approaches fine-tune LLMs on synthetic datasets annotated with uncertainty estimates. Due to the lack of ground truth confidence scores, current methods typically rely on heuristically generated proxy scores as targets, such as the model’s average accuracy over a group of similar questions (teaching, ), consistency across multiple responses (sayself, ), or model judgment (lacie, ). However, using group-level statistics as a proxy for single-instance confidence relies on the strong assumption that the questions within each group are equivalent, sampling-based methods increase both computational costs and random noise, and model judgment introduces model bias. Consequently, there remains a need for more principled and efficient approaches that more directly align an LLM’s verbalized confidence with the actual reliability of its responses.
@@
Motivated by this gap, we pose the central research question: Can LLMs be naturally calibrated during training without relying on ground-truth confidence scores or proxy confidence estimates? Our approach is inspired by the fact that classical machine learning classifiers naturally become well-calibrated during training when optimized with loss functions that are proper scoring rules (calibration1, ; calibration2, ), such as the Brier score (brier, ), which theoretically encourage the model to make probability estimates that reflect the true likelihood of correctness. Building on this insight, we introduce the notion of proper scoring rules for verbalized confidence, which formalizes the notion of a loss function that encourages LLMs to generate tokens that verbally express the true likelihood of correctness.
@@
ConfTuner provides more accurate confidence scores than the best baseline (up to 54.7% improvement in ECE and 14.4% in AUROC), and generalizes better across unseen datasets with diverse reasoning tasks, different formats of confidence expression, and even implicit confidence expressions. We also assess its effectiveness in calibrating the outputs of black-box models such as GPT-4o (gpt-4o, ). ConfTuner’s strong empirical performance suggests a meaningful alignment between its verbalized confidence and the underlying uncertainty. Beyond standard calibration metrics, we explore its broader utility in enhancing the trustworthiness of LLM-based systems. In particular, we show that well-calibrated confidence enables practical benefits, including improved LLM self-correction and better model cascade. These findings indicate that accurate confidence estimation not only enhances model interpretability and downstream performance, but also holds strong promise for advancing reliable and collaborative human–AI interaction.
@@ 2 Background: Calibration in Classification Settings
A key motivation behind our work is the intuition that binary classifiers trained using Brier score naturally become calibrated during training, without needing any extra supervision about their confidence (calibration1, ; calibration2, ). For example, when a binary classifier outputs a probability of $0.8$, we often interpret this as predicting with 80% confidence that the true label is $1$. We can do this because the classifier is trained using losses that are proper scoring rules (calibration1, ), such as Brier score. Intuitively, this means that such losses incentivize the classifier to output probabilities that reflect the model’s true likelihood of correctness. Next, we more formally define the notion of proper scoring rules.
@@
In theory, a proper scoring rule encourages the model to make probability estimates that reflect the true likelihood of correctness (calibration1, ). In particular, the Brier score $\ell_{\mathrm{B}}(p,y)=(y-p)^{2}$ has been proven to be a proper scoring rule (calibration1, ).
@@ 4.1 Experimental Setup
Datasets. Following (sayself, ), we use HotpotQA (hotpotqa, ) for training, which typically requires multi-step reasoning to derive the answer. For evaluation, besides the evaluation set of HotpotQA, we also adopt: 1) TriviaQA (triviaqa, ), which includes open-domain trivia questions and source documents; following (lacie, ), we sample 1,000 for evaluation. 2) StrategyQA (strategyqa, ), where the required reasoning steps are implicit in the question, and should be inferred strategically. 3) GSM8K (gsm8k, ), a benchmark comprising linguistically diverse and high-quality mathematics questions designed for grade school students. Here we sample 1,000 for evaluation. 4) TruthfulQA (truthfulqa, ), which evaluates how models balance factual accuracy against response utility, using questions that commonly mislead humans.
Baselines. We evaluate ConfTuner on top of three base LLMs: Llama-3.1-8B-Instruct (llama, ), Qwen2.5-7B-Instruct (qwen, ), Ministral-8B-Instruct-2410 (mistral, ) (An enhanced variant of Mistral-7B-Instruct-v0.3). For brevity, we refer to these models as LLaMA, Qwen, and Ministral, respectively, throughout the paper. We compare ConfTuner against the following baselines: 1) Base: The original, unmodified LLM. 2) Ensemble: The LLM is prompted three times to generate top-k answers with confidence, and the verbalized confidence scores are averaged to produce the final confidence estimate. 3) Two training-based methods: SaySelf (sayself, ) and LACIE (lacie, ). For LACIE, we constructed training datasets following their original implementations. For SaySelf, we directly use their training dataset (constructed based on HotpotQA). We ensure fair comparison by: i) using the same inference-time prompting strategy, and ii) re-training SaySelf and LACIE using the same base LLMs on HotpotQA. For inference, we use greedy decoding for all the methods, except for Ensemble, which requires sampling multiple responses.
Evaluation Metrics. To assess the quality of confidence estimates, we employ two metrics following previous works (canllm, ; llmtaught, ; sayself, ; lacie, ): Expected Calibration Error (ECE) (ece, ) and Area Under the ROC Curve (AUROC) (auroc, ). ECE measures the gap between a model’s predicted confidence and its empirical accuracy across probability bins, e.g., a perfectly calibrated model would achieve 80% accuracy for all samples predicted with 80% confidence. Lower ECE indicates better calibration.
@@ 4.2 Can ConfTuner Learn Effective Verbalized Confidence Estimation Capabilities?
To investigate whether ConfTuner shows good performance for verbalized confidence estimation, we conduct experiments to assess its generalization across novel datasets, different forms of confidence representation, and its adaptation to black-box models.
Generalization to Unseen Datasets. To assess ConfTuner’s generalization, we evaluate its performance on the in-distribution dataset HotpotQA (hotpotqa, ) and four out-of-distribution datasets: GSM8K (gsm8k, ), TriviaQA (triviaqa, ), StrategyQA (strategyqa, ), and TruthfulQA (truthfulqa, ). As shown in Tables  LABEL:tab:combined_ece and  LABEL:tab:combined_auroc, ConfTuner consistently achieves higher AUROC and lower ECE values across all three base models, indicating its robust generalization. Overall, training-based methods, SaySelf and LACIE, outperform the prompt-based method, Ensemble. This is primarily because even though Ensemble utilizes multiple sampling strategies, the model inherently lacks the capacity to provide reliable confidence estimates. We also illustrate ConfTuner’s accuracy among different confidence levels in Figure  3, where ConfTuner shows minimal accuracy-confidence gaps (red bars). Accuracy results and comparison to the logit-based method can be found in Appendix  F.
@@ 4.3 Can ConfTuner Help Build More Reliable and Cost-Effective LLM Systems?
ConfTuner Improves the Self-correction Ability of LLM. Self-correction offers a straightforward method to enhance LLM reliability by directly instructing the model to refine its answers (self-improvement, ). We conduct self-correction experiments on HotpotQA and TruthfulQA, where LLMs demonstrate high error rates and low confidence. Specifically, we first instruct LLM to generate answers and confidences, then retain initial responses with high confident (larger than 0.5) answers, and instruct LLM to refine low-confident (smaller than 0.5) answers. As presented in Figure  5, ConfTuner (based on Qwen) achieves larger improvements on both datasets. In contrast, baselines show marginal gains or even degradation. This is because baselines are more likely to provide low confidence for correct answers, misleading LLMs to modify correct responses into incorrect ones. The detailed accuracy results can be found in Appendix  F.
@@
ConfTuner Achieves Higher Performance Gain at Same Cost in Confidence-Based Model Cascade Systems. One important application of accurate confidence estimation is in confidence-based model cascades, where a base model’s low-confidence outputs trigger selective intervention by a stronger model to improve reliability while keeping the overall cost low. We evaluate whether the confidence estimates produced by ConfTuner can better support this process. Specifically, we compare LLaMA and its fine-tuned version, ConfTuner, by using their confidence scores to select 100 to 400 low-confidence samples for further refinement by GPT-4o (gpt-4o, ). As shown in Figure 5, ConfTuner consistently achieves higher refined accuracy, with improvements of up to 9.3% on HotpotQA and 5.5% on TruthfulQA under the same revision budget. These results show that ConfTuner’s more reliable confidence estimates enable more effective and cost-efficient cascading, improving system reliability while minimizing unnecessary interventions
@@ 5 Related Work
LLMs often struggle to reliably express their confidence (canllm, ; just, ; llmtaught, ), which may mislead users into over-relying on incorrect outputs and cause harm. Prior works (llmtaught, ; llmknow, ; Amos, ) have explored calibrating confidence scores based on the logits of LLM-generated answers, but these logits are often inaccessible to users, hindering practical use. To address this, recent studies (canllm, ; just, ; lacie, ; sayself, ; teaching, ) have focused on eliciting verbalized confidence directly from LLM outputs. Initial approaches (canllm, ; just, ) leveraged prompt strategies to guide LLMs to directly output confidence levels. While flexible, these methods often yield poorly calibrated verbalized confidence. Recent efforts (teaching, ) have shifted toward fine-tuning LLMs to produce verbalized confidence scores, typically by training models to map entire question categories to predefined confidence values. However, this category-level calibration assumes the same uncertainty scores across all questions within a class, an unrealistic premise that ignores question-level variations in difficulty or ambiguity. To overcome this, SaySelf (sayself, ) proposes question-level calibration, where confidence is estimated for individual questions. Yet, it often requires sampling multiple responses per question to infer confidence levels, which is suboptimal and incurs significant computational costs. LACIE (lacie, ) utilizes a preference dataset where responses are labeled for confidence levels. Its training objective is to encourage models to produce correct and confident or incorrect and unconfident responses. However, a key limitation of this approach is its reliance on model judgment for the initial confident/unconfident labeling, which is not accurate.
More related work for traditional calibration methods can be found in Appendix  A

### v2

@@ 1 Introduction
A large language model’s (LLM) ability to recognize and communicate uncertainty through verbalized confidence–that is, expressions of confidence conveyed in natural language, such as “I am 80 percent confident that…” (teaching)–is central to effective human–AI collaboration (li2025as). This capability is particularly important in high-stakes domains such as scientific inquiry (asai2024openscholar), law (li2024legal), and healthcare (li2024mediq), where decision quality and interpretability are essential. However, current LLMs are not explicitly trained to express calibrated uncertainty. As a result, they often report very high confidence even when producing hallucinated or incorrect content (hallucination1; hallucination2; hallucination3; canllm). This overconfidence problem undermines trust and poses serious challenges for the safe deployment of LLMs (Figure 1).
Recent efforts (just; canllm; teaching; sayself; lacie) have focused on improving the elicitation of verbalized confidence from LLMs. Prompt-based methods rely on carefully crafted instructions (just; canllm), but have shown limited effects in improving calibration (just; canllm). Alternatively, training-based approaches fine-tune LLMs on synthetic datasets annotated with uncertainty estimates. Due to the lack of ground truth confidence scores, current methods typically rely on heuristically generated proxy scores as targets, such as the model’s average accuracy over a group of similar questions (teaching), consistency across multiple responses (sayself), or model judgment (lacie). However, using group-level statistics as a proxy for single-instance confidence relies on the strong assumption that the questions within each group are equivalent, sampling-based methods increase both computational costs and random noise, and model judgment introduces model bias. Consequently, there remains a need for more principled and efficient approaches that more directly align an LLM’s verbalized confidence with the actual reliability of its responses.
@@
Motivated by this gap, we pose the central research question: Can LLMs be naturally calibrated during training without relying on ground-truth confidence scores or proxy confidence estimates? Our approach is inspired by the fact that classical machine learning classifiers naturally become well-calibrated during training when optimized with loss functions that are proper scoring rules (calibration1; calibration2), such as the Brier score (brier), which theoretically encourage the model to make probability estimates that reflect the true likelihood of correctness. Building on this insight, we introduce the notion of proper scoring rules for verbalized confidence, which formalizes the notion of a loss function that encourages LLMs to generate tokens that verbally express the true likelihood of correctness.
@@
ConfTuner provides more accurate confidence scores than the best baseline (up to 54.7% improvement in ECE and 14.4% in AUROC), and generalizes better across unseen datasets with diverse reasoning tasks, different formats of confidence expression, and even implicit confidence expressions. We also assess its effectiveness in calibrating the outputs of black-box models such as GPT-4o (gpt-4o). ConfTuner’s strong empirical performance suggests a meaningful alignment between its verbalized confidence and the underlying uncertainty. Beyond standard calibration metrics, we explore its broader utility in enhancing the trustworthiness of LLM-based systems. In particular, we show that well-calibrated confidence enables practical benefits, including improved LLM self-correction and better model cascade. These findings indicate that accurate confidence estimation not only enhances model interpretability and downstream performance, but also holds strong promise for advancing reliable and collaborative human–AI interaction.
@@ 2 Background: Calibration in Classification Settings
A key motivation behind our work is the intuition that binary classifiers trained using Brier score naturally become calibrated during training, without needing any extra supervision about their confidence (calibration1; calibration2). For example, when a binary classifier outputs a probability of $0.8$, we often interpret this as predicting with 80% confidence that the true label is $1$. We can do this because the classifier is trained using losses that are proper scoring rules (calibration1), such as Brier score. Intuitively, this means that such losses incentivize the classifier to output probabilities that reflect the model’s true likelihood of correctness. Next, we more formally define the notion of proper scoring rules.
@@
In theory, a proper scoring rule encourages the model to make probability estimates that reflect the true likelihood of correctness (calibration1). In particular, the Brier score $\ell_{\mathrm{B}}(p,y)=(y-p)^{2}$ has been proven to be a proper scoring rule (calibration1).
@@ 4.1 Experimental Setup
Datasets. Following (sayself), we use HotpotQA (hotpotqa) for training, which typically requires multi-step reasoning to derive the answer. For evaluation, besides the evaluation set of HotpotQA, we also adopt: 1) TriviaQA (triviaqa), which includes open-domain trivia questions and source documents; following (lacie), we sample 1,000 for evaluation. 2) StrategyQA (strategyqa), where the required reasoning steps are implicit in the question, and should be inferred strategically. 3) GSM8K (gsm8k), a benchmark comprising linguistically diverse and high-quality mathematics questions designed for grade school students. Here we sample 1,000 for evaluation. 4) TruthfulQA (truthfulqa), which evaluates how models balance factual accuracy against response utility, using questions that commonly mislead humans.
Baselines. We evaluate ConfTuner on top of three base LLMs: Llama-3.1-8B-Instruct (llama), Qwen2.5-7B-Instruct (qwen), Ministral-8B-Instruct-2410 (mistral) (An enhanced variant of Mistral-7B-Instruct-v0.3). For brevity, we refer to these models as LLaMA, Qwen, and Ministral, respectively, throughout the paper. We compare ConfTuner against the following baselines: 1) Base: The original, unmodified LLM. 2) Ensemble: The LLM is prompted three times to generate top-k answers with confidence, and the verbalized confidence scores are averaged to produce the final confidence estimate. 3) Two training-based methods: SaySelf (sayself) and LACIE (lacie). For LACIE, we constructed training datasets following their original implementations. For SaySelf, we directly use their training dataset (constructed based on HotpotQA). We ensure fair comparison by: i) using the same inference-time prompting strategy, and ii) re-training SaySelf and LACIE using the same base LLMs on HotpotQA. For inference, we use greedy decoding for all the methods, except for Ensemble, which requires sampling multiple responses.
Evaluation Metrics. To assess the quality of confidence estimates, we employ two metrics following previous works (canllm; llmtaught; sayself; lacie): Expected Calibration Error (ECE) (ece) and Area Under the ROC Curve (AUROC) (auroc). ECE measures the gap between a model’s predicted confidence and its empirical accuracy across probability bins, e.g., a perfectly calibrated model would achieve 80% accuracy for all samples predicted with 80% confidence. Lower ECE indicates better calibration.
@@ 4.2 Can ConfTuner Learn Effective Verbalized Confidence Estimation Capabilities?
To investigate whether ConfTuner shows good performance for verbalized confidence estimation, we conduct experiments to assess its generalization across novel datasets, different forms of confidence representation, implicit confidence expressions, and its adaptation to black-box models.
Generalization to Unseen Datasets. To assess ConfTuner’s generalization, we evaluate its performance on the in-distribution dataset HotpotQA (hotpotqa) and four out-of-distribution datasets: GSM8K (gsm8k), TriviaQA (triviaqa), StrategyQA (strategyqa), and TruthfulQA (truthfulqa). As shown in Tables  LABEL:tab:combined_ece and  LABEL:tab:combined_auroc, ConfTuner consistently achieves higher AUROC and lower ECE values across all three base models, indicating its robust generalization. Overall, training-based methods, SaySelf and LACIE, outperform the prompt-based method, Ensemble. This is primarily because even though Ensemble utilizes multiple sampling strategies, the model inherently lacks the capacity to provide reliable confidence estimates. We also illustrate ConfTuner’s accuracy among different confidence levels in Figure  3, where ConfTuner shows minimal accuracy-confidence gaps (red bars). Accuracy results and comparison to the logit-based method can be found in Appendix  F.
@@ 4.3 Can ConfTuner Help Build More Reliable and Cost-Effective LLM Systems?
ConfTuner Improves the Self-correction Ability of LLM. Self-correction offers a straightforward method to enhance LLM reliability by directly instructing the model to refine its answers (self-improvement). We conduct self-correction experiments on HotpotQA and TruthfulQA, where LLMs demonstrate high error rates and low confidence. Specifically, we first instruct LLM to generate answers and confidences, then retain initial responses with high confident (larger than 0.5) answers, and instruct LLM to refine low-confident (smaller than 0.5) answers. As presented in Figure  5, ConfTuner (based on Qwen) achieves larger improvements on both datasets. In contrast, baselines show marginal gains or even degradation. This is because baselines are more likely to provide low confidence for correct answers, misleading LLMs to modify correct responses into incorrect ones. The detailed accuracy results can be found in Appendix  F.
@@
ConfTuner Achieves Higher Performance Gain at Same Cost in Confidence-Based Model Cascade Systems. One important application of accurate confidence estimation is in confidence-based model cascades, where a base model’s low-confidence outputs trigger selective intervention by a stronger model to improve reliability while keeping the overall cost low. We evaluate whether the confidence estimates produced by ConfTuner can better support this process. Specifically, we compare LLaMA and its fine-tuned version, ConfTuner, by using their confidence scores to select 100 to 400 low-confidence samples for further refinement by GPT-4o (gpt-4o). As shown in Figure 5, ConfTuner consistently achieves higher refined accuracy, with improvements of up to 9.3% on HotpotQA and 5.5% on TruthfulQA under the same revision budget. These results show that ConfTuner’s more reliable confidence estimates enable more effective and cost-efficient cascading, improving system reliability while minimizing unnecessary interventions
@@ 5 Related Work
LLMs often struggle to reliably express their confidence (canllm; just; llmtaught), which may mislead users into over-relying on incorrect outputs and cause harm. Prior works (llmtaught; llmknow; Amos) have explored calibrating confidence scores based on the logits of LLM-generated answers, but these logits are often inaccessible to users, hindering practical use. To address this, recent studies (canllm; just; lacie; sayself; teaching) have focused on eliciting verbalized confidence directly from LLM outputs. Initial approaches (canllm; just) leveraged prompt strategies to guide LLMs to directly output confidence levels. While flexible, these methods often yield poorly calibrated verbalized confidence. Recent efforts (teaching) have shifted toward fine-tuning LLMs to produce verbalized confidence scores, typically by training models to map entire question categories to predefined confidence values. However, this category-level calibration assumes the same uncertainty scores across all questions within a class, an unrealistic premise that ignores question-level variations in difficulty or ambiguity. To overcome this, SaySelf (sayself) proposes question-level calibration, where confidence is estimated for individual questions. Yet, it often requires sampling multiple responses per question to infer confidence levels, which is suboptimal and incurs significant computational costs. LACIE (lacie) utilizes a preference dataset where responses are labeled for confidence levels. Its training objective is to encourage models to produce correct and confident or incorrect and unconfident responses. However, a key limitation of this approach is its reliance on model judgment for the initial confident/unconfident labeling, which is not accurate.
More related work for traditional calibration methods can be found in Appendix  A.